### v4

@@
[…]
@@ Introduction
$\xi^{0}_{\star}$ $=\mathbb{V}\mathrm{a}\mathbb{R}_{\alpha}(X_{0})\coloneqq\inf{\big\{\xi\in\mathbb{R}\colon\mathbb{P}(X_{0}\leq\xi)\geq\alpha\big\}},$
$\chi^{0}_{\star}$ $=\mathbb{E}\mathbb{S}_{\alpha}(X_{0})\coloneqq\frac{1}{1-\alpha}\int_{\alpha}^{1}\mathbb{V}\mathrm{a}\mathbb{R}_{a}(X_{0})\mathrm{d}a=\mathbb{E}[X_{0}\,|\,X_{0}\geq\xi^{0}_{\star}]=\frac{1}{1-\alpha}\mathbb{E}[X_{0}\mathds{1}_{X_{0}\geq\xi^{0}_{\star}}].$ (1)
@@
$\min_{\xi\in\mathbb{R}}{V_{0}(\xi)},\quad\mbox{ where }\quad V_{0}(\xi)\coloneqq\xi+\frac{1}{1-\alpha}\mathbb{E}[(X_{0}-\xi)^{+}].$ (2)
@@
$X_{0}=\mathbb{E}[\varphi(Y,Z)|Y],$ (3)
where $Y\in\mathbb{R}^{d}$ and $Z\in\mathbb{R}^{q}$ are independent random variables and $\varphi\colon\mathbb{R}^{d}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is a measurable function such that $\varphi(Y,Z)\in L^{1}(\mathbb{P})$. In financial applications, $\varphi(Y,Z)$ models the future cash flows of a portfolio, $Y$ the underlying risk factors up to a certain time horizon, $Z$ the underlying risk factors beyond that time horizon, and $X_{0}$ the loss of the portfolio at that time horizon.
@@
The paper is structured as follows. Sections 14 introduce the NSA, averaged NSA (ANSA), MLSA and averaged MLSA (AMLSA) algorithms, establish CLTs for their renormalized VaR and ES estimation errors (in Corollaries 1.4 and 2.4 and Theorems 3.5 and 4.1) and discuss their complexities and related parametrizations (in Thoerems 1.8, 2.8, 3.8 and 4.4). Section 5 presents a numerical case study supporting the paper’s theoretical findings.
@@ 1 Nested Stochastic Approximation Algorithm
The brute force, naively nested SA algorithm introduced in [11] for solving (2)–(3) is based on defining the approximate problem
$\min_{\xi\in\mathbb{R}}{V_{h}(\xi)},\quad\mbox{ where }\quad V_{h}(\xi)\coloneqq\xi+\frac{1}{1-\alpha}\mathbb{E}[(X_{h}-\xi)^{+}]$ (4)
@@
and $\{Z^{(k)},1\leq k\leq K\}\mathrel{{\mathop{\sim}\limits^{\text{\tiny\rm i.i.d.}}}}Z$ are independent of $Y$. $X_{h}$, $h\in\mathcal{H}$, represents a biased (Monte Carlo) estimator of $X_{0}$ that, in principle, converges -$\mathbb{P}\text{\rm-as}$ to $X_{0}$ under suitable assumptions via the conditional law of large numbers. Note that the notion of bias here is linked to the ensuing solution $(\xi^{h}_{\star},\chi^{h}_{\star})$ to (4). Indeed, for $h\in\mathcal{H}$, if the cdf $F_{X_{h}}$ of $X_{h}$ is continuous and increasing, then
$\xi^{h}_{\star}=\operatorname*{argmin}{V_{h}}\quad\mbox{ and }\quad\chi^{h}_{\star}=\min{V_{h}}$ (5)
@@
$V_{h}^{\prime}(\xi)=\frac{1}{1-\alpha}(F_{X_{h}}(\xi)-\alpha)=\mathbb{E}[H_{1}(\xi,X_{h})],\quad H_{1}(\xi,x)=1-\frac{1}{1-\alpha}\mathds{1}_{x\geq\xi},\quad\xi,x\in\mathbb{R}.$ (6)
@@
In view of (6), if $F_{X_{h}}$ is continuous and increasing, then $V_{h}$ is strictly convex so that $\xi^{h}_{\star}$ is unique and $\left\{\xi^{h}_{\star}\right\}=\left\{V^{\prime}_{h}=0\right\}$. This hints at a stochastic gradient descent type algorithm based on (6) to approximate $\xi^{h}_{\star}$. Assuming that such an algorithm is driven by innovations $\{X_{h}^{(n)},n\geq 1\}\mathrel{{\mathop{\sim}\limits^{\text{\tiny\rm i.i.d.}}}}X_{h}$ and produces iterates $\{\xi^{h}_{n},n\geq 0\}$ converging to $\xi^{h}_{\star}$ -$\mathbb{P}\text{\rm-as}$, then (4), (5), the law of large numbers together with Cesàro’s lemma yield under some suitable assumptions:
$\chi^{h}_{\star}=\mathbb{E}\Big[\xi^{h}_{\star}+\frac{1}{1-\alpha}(X_{h}-\xi^{h}_{\star})^{+}\Big]=\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{k=1}^{n}\Big(\xi^{h}_{k-1}+\frac{1}{1-\alpha}(X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\Big).$
@@ 1.1 Convergence Rate Analysis
We postulate below some essential assumptions on $\{X_{h},h\in\mathcal{H}\cup\{0\}\}$. {assumption}[[11, Assumptions 2.2 & 2.5]]
@@
    for some functions $v,w(\,\cdot\,,h)\colon\mathbb{R}\rightarrow\mathbb{R}$ satisfying, for any $\xi^{0}_{\star}\in\operatorname*{Argmin}{V_{0}}$,
@@
    ([11, Remarks 2.3 & 2.6]).  Assumptions 1.1(i) and 1.1(ii) are guaranteed under the conditions of [24, Propositions 5.1(a) & 5.1(b)]. The last part of Assumption 1.1(i) reads $w(\xi^{0}_{\star},h)=\int_{\xi^{0}_{\star}}^{\infty}w(\xi,h)\mathrm{d}\xi=\mathrm{o}(1)$ as $\mathcal{H}\ni h\downarrow 0$, $\xi^{0}_{\star}\in\operatorname*{Argmin}{V_{0}}$.
    Assumption 1.1(iii) is natural in view of Assumption 1.1(ii) and an increasing $F_{X_{0}}$. It may be weakened to $f_{X_{0}}(\xi^{0}_{\star})>0$, but the current formulation eases the obtention of the Lyapunov properties [11, Lemma D.1] that play a central role in controlling the statistical error [11, Theorem 2.7] (Lemma A.3(ii), especially the constant defined in (9) whose value must be positive).
@@
The global error of the NSA scheme (7)-(8) writes as a sum of statistical and bias errors:
@@
$\lambda=\inf_{h\in\mathcal{H}\cup\{0\}}\frac{3}{8}V_{h}^{\prime\prime}(\xi^{h}_{\star})\wedge\|V_{h}^{\prime\prime}\|_{\infty}\frac{V_{h}^{\prime\prime}(\xi^{h}_{\star})^{4}}{4[V_{h}^{\prime\prime}]_{\mathrm{Lip}}^{2}}>0.$ (9)
@@
Suppose that Assumptions 1.1 and 1.1 hold, and that $\mathbb{E}[|\varphi(Y,Z)|^{2+\delta}]<\infty$ for some $\delta>0$. If $\gamma_{n}=\gamma_{1}n^{-\beta}$, $n\geq 1$, $\beta\in\big(\frac{1}{2},1\big]$, with $\lambda\gamma_{1}>1$ if $\beta=1$, then
$\begin{pmatrix}h^{-\beta}\big(\xi^{h}_{\lceil h^{-2}\rceil}-\xi^{h}_{\star}\big)\\ h^{-1}\big(\chi^{h}_{\lceil h^{-2}\rceil}-\chi^{h}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\Sigma_{\beta})\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
@@
$\Sigma_{\beta}=\begin{pmatrix}\frac{\alpha\gamma_{1}}{2f_{X_{0}}(\xi^{0}_{\star})-\gamma_{1}^{-1}(1-\alpha)\mathds{1}_{\beta=1}}&\alpha\frac{\chi^{0}_{\star}-\xi^{0}_{\star}}{f_{X_{0}}(\xi^{0}_{\star})}\mathds{1}_{\beta=1}\\ \alpha\frac{\chi^{0}_{\star}-\xi^{0}_{\star}}{f_{X_{0}}(\xi^{0}_{\star})}\mathds{1}_{\beta=1}&\frac{\mathbb{V}\mathrm{ar}((X_{0}-\xi^{0}_{\star})^{+})}{(1-\alpha)^{2}}\end{pmatrix}$. (10)
Proof 1.3.
See Appendix B.
@@
$\begin{pmatrix}h^{-\beta}\big(\xi^{h}_{\lceil h^{-2}\rceil}-\xi^{0}_{\star}\big)\\ h^{-1}\big(\chi^{h}_{\lceil h^{-2}\rceil}-\chi^{0}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\left(\begin{pmatrix}-\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\mathds{1}_{\beta=1}\\ -\int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix},\Sigma_{\beta}\right)\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
where $\Sigma_{\beta}$ is given by (10).
Proof 1.5.
@@
$\begin{pmatrix}h^{-\beta}(\xi^{h}_{\star}-\xi^{0}_{\star})\\ h^{-1}(\chi^{h}_{\star}-\chi^{0}_{\star})\end{pmatrix}\rightarrow-\begin{pmatrix}\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\mathds{1}_{\beta=1}\\ \int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
Combining this limit with Theorem 1.2 yields the desired result.
@@
    While the ES converges in $\mathrm{O}(h)$ for any $\beta\in\big(\frac{1}{2},1\big]$, the VaR converges in $\mathrm{O}(h^{\beta})$, only reaching optimality at $\beta=1$. This comes at the cost of an additional non trivial constraint on $\gamma_{1}$ stipulating that $\lambda\gamma_{1}>1$, the value of $\lambda$ being inaccessible in view of (9). Such restriction is classical in the stochastic approximation literature [13].
@@
    The VaR error’s variance factor in (10) is affected by the choice of $\gamma_{1}$.
    When $\beta\in\big(\frac{1}{2},1\big)$, this factor is linear in $\gamma_{1}$, suggesting to take $\gamma_{1}$ small. But, by (7), $\lim_{\gamma_{1}\downarrow 0}\xi^{h}_{n}=\xi^{h}_{0}$, i.e. the algorithm does not learn for too small a value of $\gamma_{1}$. A fine-tuning of $\gamma_{1}$ must therefore be carried out.
    The optimal setting $\beta=1$ affects this factor differently. While it evolves in $\mathrm{O}(\gamma_{1})$ for $\gamma_{1}\uparrow\infty$, it evolves in $\mathrm{O}\big(\big(\gamma_{1}-\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}\big)^{-1}\big)$ when $\gamma_{1}\downarrow\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}$, reflecting a high numerical instability. Taking $\gamma_{1}=\frac{1-\alpha}{f_{X_{0}}(\xi^{0}_{\star})}$ to set this factor to its minimum $\frac{\alpha(1-\alpha)}{f_{X_{0}}(\xi^{0}_{\star})^{2}}$ is infeasible as neither $f_{X_{0}}$ nor $\xi^{0}_{\star}$ are a priori accessible. Hence an upstream fine-tuning of $\gamma_{1}$ is required.
@@
    The ES error’s variance factor is independent of $\gamma_{1}$, reflecting a stronger numerical stability for the ES estimation. This is to be nuanced, as ES dependency on $\gamma_{1}$ is dissimulated in secondary error terms in (42) that vanish faster than the main martingale component giving the CLT 1.4.
@@
    Although analytic, the covariance matrix (10) is not computable, as it depends on $f_{X_{0}}$, $\xi^{0}_{\star}$ and $\chi^{0}_{\star}$ that are a priori unknown. It must therefore be statistically estimated using several samplings of the renormalized estimation errors, in order to get confidence regions. Alternatively, as in [8, Remark 3.4], the ES variance factor can be estimated using the Monte Carlo approximation
    $\mathbb{V}\mathrm{ar}\big((X_{0}-\xi^{0}_{\star})^{+}\big)\approx\frac{1}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\big((X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\big)^{2}-\Big(\frac{1}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}(X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\Big)^{2}.$ (11)
@@
    [9, Theorem 1.3] relies on [29, Theorem 1] to analyze a two-time-scale VaR-ES unbiased scheme. Unlike (8), their ES recursion writes
    $\chi^{0}_{n+1}=(1-b_{n+1})\chi^{0}_{n}+b_{n+1}X_{0}^{(n+1)}\mathds{1}_{X_{0}^{(n+1)}>\xi^{0}_{n}},$ (12)
    echoing the ES definition (1). The $\{\xi^{0}_{k},k\geq 1\}$ are supposingly iterates produced by (7) by swapping $X_{h}$ with an exact simulation of $X_{0}$ and the $\{b_{n},n\geq 1\}$ are damping weights.
@@
    [18, Theorem 2.7] presents a biased CLT similar to ours, however, it relies on the mean-reverting property [18, Assumption (HMR)], stipulating that the objective function being minimized is strongly convex. This is not the case with (4). Moreover, the aforementioned result concerns only the VaR component of our NSA scheme.

### v5

@@
Abstract. Crépey, Frikha, and Louzi (2025) introduced a nested stochastic approximation algorithm and its multilevel acceleration to compute the value-at-risk and expected shortfall of a random financial loss. We hereby establish central limit theorems for the renormalized estimation errors associated with both algorithms as well as their averaged versions. Our findings are substantiated through a numerical example.
Keywords. value-at-risk ; expected shortfall ; stochastic approximation ; multilevel Monte Carlo ; Polyak-Ruppert averaging ; central limit theorem
MSC. 65C05 ; 62L20 ; 62G32 ; 91Gxx
DOI. 10.1214/24-EJP1246    arXiv. 2311.15333    HAL. 04304985
@@ Introduction
$\xi^{0}_{\star}$ $=\operatorname{VaR}_{\alpha}(X_{0})\coloneqq\inf{\big\{\xi\in\mathbb{R}\colon\mathbb{P}(X_{0}\leq\xi)\geq\alpha\big\}},$
$\chi^{0}_{\star}$ $=\operatorname{ES}_{\alpha}(X_{0})\coloneqq\frac{1}{1-\alpha}\int_{\alpha}^{1}\operatorname{VaR}_{a}(X_{0})\mathrm{d}a=\mathbb{E}[X_{0}\,|\,X_{0}\geq\xi^{0}_{\star}]=\frac{1}{1-\alpha}\mathbb{E}[X_{0}\mathds{1}_{X_{0}\geq\xi^{0}_{\star}}].$ (0.1)
@@
$\min_{\xi\in\mathbb{R}}{V_{0}(\xi)},\quad\mbox{ where }\quad V_{0}(\xi)\coloneqq\xi+\frac{1}{1-\alpha}\mathbb{E}[(X_{0}-\xi)^{+}].$ (0.2)
@@
$X_{0}=\mathbb{E}[\varphi(Y,Z)|Y],$ (0.3)
where $Y\in\mathbb{R}^{d}$ and $Z\in\mathbb{R}^{q}$ are independent random variables and $\varphi\colon\mathbb{R}^{d}\times\mathbb{R}^{q}\to\mathbb{R}$ is a measurable function such that $\varphi(Y,Z)\in L^{1}(\mathbb{P})$. In financial applications, $\varphi(Y,Z)$ models the future cash flows of a portfolio, $Y$ the underlying risk factors up to a certain time horizon, $Z$ the underlying risk factors beyond that time horizon, and $X_{0}$ the loss of the portfolio at that time horizon.
@@
The paper is structured as follows. Sections 14 introduce the NSA, averaged NSA (ANSA), MLSA and averaged MLSA (AMLSA) algorithms, establish CLTs for their renormalized VaR and ES estimation errors (in Corollaries 1.5 and 2.3 and Theorems 3.5 and 4.1) and discuss their complexities and related parametrizations (in Thoerems 1.8, 2.6, 3.7 and 4.3). Section 5 presents a numerical case study supporting the paper’s theoretical findings.
@@ 1 Nested Stochastic Approximation Algorithm
The brute force, naively nested SA algorithm introduced in [11] for solving (0.2)–(0.3) is based on defining the approximate problem
$\min_{\xi\in\mathbb{R}}{V_{h}(\xi)},\quad\mbox{ where }\quad V_{h}(\xi)\coloneqq\xi+\frac{1}{1-\alpha}\mathbb{E}[(X_{h}-\xi)^{+}]$ (1.1)
@@
and $\{Z^{(k)},1\leq k\leq K\}\mathrel{{\mathop{\sim}\limits^{\text{\tiny\rm i.i.d.}}}}Z$ are independent of $Y$. $X_{h}$, $h\in\mathcal{H}$, represents a biased (Monte Carlo) estimator of $X_{0}$ that, in principle, converges -$\operatorname{\mathds{P}\text{-as}}$ to $X_{0}$ under suitable assumptions via the conditional law of large numbers. Note that the notion of bias here is linked to the ensuing solution $(\xi^{h}_{\star},\chi^{h}_{\star})$ to (1.1). Indeed, for $h\in\mathcal{H}$, if the cdf $F_{X_{h}}$ of $X_{h}$ is continuous and increasing, then
$\xi^{h}_{\star}=\operatorname*{argmin}{V_{h}}\quad\mbox{ and }\quad\chi^{h}_{\star}=\min{V_{h}}$ (1.2)
@@
$V_{h}^{\prime}(\xi)=\frac{1}{1-\alpha}(F_{X_{h}}(\xi)-\alpha)=\mathbb{E}[H_{1}(\xi,X_{h})],\quad H_{1}(\xi,x)=1-\frac{1}{1-\alpha}\mathds{1}_{x\geq\xi},\quad\xi,x\in\mathbb{R}.$ (1.3)
@@
In view of (1.3), if $F_{X_{h}}$ is continuous and increasing, then $V_{h}$ is strictly convex so that $\xi^{h}_{\star}$ is unique and $\left\{\xi^{h}_{\star}\right\}=\left\{V^{\prime}_{h}=0\right\}$. This hints at a stochastic gradient descent type algorithm based on (1.3) to approximate $\xi^{h}_{\star}$. Assuming that such an algorithm is driven by innovations $\{X_{h}^{(n)},n\geq 1\}\mathrel{{\mathop{\sim}\limits^{\text{\tiny\rm i.i.d.}}}}X_{h}$ and produces iterates $\{\xi^{h}_{n},n\geq 0\}$ converging to $\xi^{h}_{\star}$ -$\operatorname{\mathds{P}\text{-as}}$, then (1.1), (1.2), the law of large numbers together with Cesàro’s lemma yield under some suitable assumptions:
$\chi^{h}_{\star}=\mathbb{E}\Big[\xi^{h}_{\star}+\frac{1}{1-\alpha}(X_{h}-\xi^{h}_{\star})^{+}\Big]=\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\Big(\xi^{h}_{k-1}+\frac{1}{1-\alpha}(X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\Big).$
@@ 1.1 Convergence Rate Analysis
We postulate below some essential assumptions on $\{X_{h},h\in\mathcal{H}\cup\{0\}\}$.
Assumption 1.1 ([11, Assumptions 2.2 & 2.5]).
@@
    for some functions $v,w(\,\cdot\,,h)\colon\mathbb{R}\to\mathbb{R}$ satisfying, for any $\xi^{0}_{\star}\in\operatorname*{Argmin}{V_{0}}$,
@@
    ([11, Remarks 2.3 & 2.6]).  Assumptions 1.1(i) and 1.1(ii) are guaranteed under the conditions of [24, Propositions 5.1(a) & 5.1(b)]. The last part of Assumption 1.1(i) reads $w(\xi^{0}_{\star},h)=\int_{\xi^{0}_{\star}}^{\infty}w(\xi,h)\mathrm{d}\xi=\operatorname{o}(1)$ as $\mathcal{H}\ni h\downarrow 0$, $\xi^{0}_{\star}\in\operatorname*{Argmin}{V_{0}}$.
    Assumption 1.1(iii) is natural in view of Assumption 1.1(ii) and an increasing $F_{X_{0}}$. It may be weakened to $f_{X_{0}}(\xi^{0}_{\star})>0$, but the current formulation eases the obtention of the Lyapunov properties [11, Lemma D.1] that play a central role in controlling the statistical error [11, Theorem 2.7] (Lemma A.3(ii), especially the constant defined in (1.6) whose value must be positive).
@@
The global error of the NSA scheme (1.4)-(1.5) writes as a sum of statistical and bias errors:
@@
$\lambda=\inf_{h\in\mathcal{H}\cup\{0\}}\frac{3}{8}V_{h}^{\prime\prime}(\xi^{h}_{\star})\wedge\|V_{h}^{\prime\prime}\|_{\infty}\frac{V_{h}^{\prime\prime}(\xi^{h}_{\star})^{4}}{4[V_{h}^{\prime\prime}]_{\mathrm{Lip}}^{2}}>0.$ (1.6)
@@
Suppose that Assumptions 1.1 and 1.3 hold, and that $\mathbb{E}[|\varphi(Y,Z)|^{2+\delta}]<\infty$ for some $\delta>0$. If $\gamma_{n}=\gamma_{1}n^{-\beta}$, $n\geq 1$, $\beta\in\big(\frac{1}{2},1\big]$, with $\lambda\gamma_{1}>1$ if $\beta=1$, then
$\begin{pmatrix}h^{-\beta}\big(\xi^{h}_{\lceil h^{-2}\rceil}-\xi^{h}_{\star}\big)\\ h^{-1}\big(\chi^{h}_{\lceil h^{-2}\rceil}-\chi^{h}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\Sigma_{\beta})\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
@@
$\Sigma_{\beta}=\begin{pmatrix}\frac{\alpha\gamma_{1}}{2f_{X_{0}}(\xi^{0}_{\star})-\gamma_{1}^{-1}(1-\alpha)\mathds{1}_{\beta=1}}&\alpha\frac{\chi^{0}_{\star}-\xi^{0}_{\star}}{f_{X_{0}}(\xi^{0}_{\star})}\mathds{1}_{\beta=1}\\ \alpha\frac{\chi^{0}_{\star}-\xi^{0}_{\star}}{f_{X_{0}}(\xi^{0}_{\star})}\mathds{1}_{\beta=1}&\frac{\operatorname{\mathbb{V}ar}((X_{0}-\xi^{0}_{\star})^{+})}{(1-\alpha)^{2}}\end{pmatrix}$. (1.7)
Proof.
See Appendix B. ∎
@@
$\begin{pmatrix}h^{-\beta}\big(\xi^{h}_{\lceil h^{-2}\rceil}-\xi^{0}_{\star}\big)\\ h^{-1}\big(\chi^{h}_{\lceil h^{-2}\rceil}-\chi^{0}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\left(\begin{pmatrix}-\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\mathds{1}_{\beta=1}\\ -\int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix},\Sigma_{\beta}\right)\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
where $\Sigma_{\beta}$ is given by (1.7).
Proof.
@@
$\begin{pmatrix}h^{-\beta}(\xi^{h}_{\star}-\xi^{0}_{\star})\\ h^{-1}(\chi^{h}_{\star}-\chi^{0}_{\star})\end{pmatrix}\to-\begin{pmatrix}\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\mathds{1}_{\beta=1}\\ \int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
Combining this limit with Theorem 1.4 yields the desired result. ∎
@@
    While the ES converges in $\operatorname{O}(h)$ for any $\beta\in\big(\frac{1}{2},1\big]$, the VaR converges in $\operatorname{O}(h^{\beta})$, only reaching optimality at $\beta=1$. This comes at the cost of an additional non trivial constraint on $\gamma_{1}$ stipulating that $\lambda\gamma_{1}>1$, the value of $\lambda$ being inaccessible in view of (1.6). Such restriction is classical in the stochastic approximation literature [13].
@@
    The VaR error’s variance factor in (1.7) is affected by the choice of $\gamma_{1}$.
    When $\beta\in\big(\frac{1}{2},1\big)$, this factor is linear in $\gamma_{1}$, suggesting to take $\gamma_{1}$ small. But, by (1.4), $\lim_{\gamma_{1}\downarrow 0}\xi^{h}_{n}=\xi^{h}_{0}$, i.e. the algorithm does not learn for too small a value of $\gamma_{1}$. A fine-tuning of $\gamma_{1}$ must therefore be carried out.
    The optimal setting $\beta=1$ affects this factor differently. While it evolves in $\operatorname{O}(\gamma_{1})$ for $\gamma_{1}\uparrow\infty$, it evolves in $\operatorname{O}\big(\big(\gamma_{1}-\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}\big)^{-1}\big)$ when $\gamma_{1}\downarrow\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}$, reflecting a high numerical instability. Taking $\gamma_{1}=\frac{1-\alpha}{f_{X_{0}}(\xi^{0}_{\star})}$ to set this factor to its minimum $\frac{\alpha(1-\alpha)}{f_{X_{0}}(\xi^{0}_{\star})^{2}}$ is infeasible as neither $f_{X_{0}}$ nor $\xi^{0}_{\star}$ are a priori accessible. Hence an upstream fine-tuning of $\gamma_{1}$ is required.
@@
    The ES error’s variance factor is independent of $\gamma_{1}$, reflecting a stronger numerical stability for the ES estimation. This is to be nuanced, as ES dependency on $\gamma_{1}$ is dissimulated in secondary error terms in (B.8) that vanish faster than the main martingale component giving the CLT 1.5.
@@
    Although analytic, the covariance matrix (1.7) is not computable, as it depends on $f_{X_{0}}$, $\xi^{0}_{\star}$ and $\chi^{0}_{\star}$ that are a priori unknown. It must therefore be statistically estimated using several samplings of the renormalized estimation errors, in order to get confidence regions. Alternatively, as in [8, Remark 3.4], the ES variance factor can be estimated using the Monte Carlo approximation
    $\operatorname{\mathbb{V}ar}\big((X_{0}-\xi^{0}_{\star})^{+}\big)\approx\frac{1}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\big((X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\big)^{2}-\Big(\frac{1}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}(X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\Big)^{2}.$ (1.8)
@@
    [9, Theorem 1.3] relies on [29, Theorem 1] to analyze a two-time-scale VaR-ES unbiased scheme. Unlike (1.5), their ES recursion writes
    $\chi^{0}_{n+1}=(1-b_{n+1})\chi^{0}_{n}+b_{n+1}X_{0}^{(n+1)}\mathds{1}_{X_{0}^{(n+1)}>\xi^{0}_{n}},$ (1.9)
    echoing the ES definition (0.1). The $\{\xi^{0}_{k},k\geq 1\}$ are supposingly iterates produced by (1.4) by swapping $X_{h}$ with an exact simulation of $X_{0}$ and the $\{b_{n},n\geq 1\}$ are damping weights.
@@
    [18, Theorem 2.7] presents a biased CLT similar to ours, however, it relies on the mean-reverting property [18, Assumption (HMR)], stipulating that the objective function being minimized is strongly convex. This is not the case with (1.1). Moreover, the aforementioned result concerns only the VaR component of our NSA scheme.